# Monte Carlo Tree Diffusion with Multiple Experts for Protein Design

### v1

@@ 4 The MCTD-ME Algorithm
Progressive pLDDT Masking for Targeted Denoising. Rather than resampling all positions uniformly, we exploit predicted structural confidence (pLDDT) to target only low-confidence residues. Specifically, at reverse step $t$ we define a mask $\mathcal{M}_{t}$ by thresholding the pLDDT profile of $y_{t}$, masking unstable regions while preserving high-confidence subsequences. Crucially, the masking threshold decreases over diffusion steps, progressively reducing the fraction of masked residues as denoising proceeds. This “progressive masking” strategy focuses computation on the most uncertain regions early on, while stabilizing promising motifs in later stages:
@@
This masked input encourages structural refinement and prevents overwriting already stable subsequences, reducing unnecessary sampling noise and improving convergence.
Multi-Expert Guided Expansion via Rollouts. To enrich proposal diversity and better imitate the training distribution, we leverage multiple pretrained experts during expansion. Given a masked sequence $\text{Mask}(y_{t};\mathcal{M}_{t})$, each expert $e\in\{1,\dots,E\}$ performs $k$ rollouts by sampling its conditional distribution $p_{\phi}^{(e)}(\cdot\mid y_{t},\mathcal{M}_{t})$. This yields a pooled candidate set
@@
where $y_{t-1}^{(e,r)}$ denotes a complete sequence obtained by filling the masked sites and splicing them back into $y_{t}$. All candidates in $\mathcal{C}_{t}$ are de-duplicated and scored once with the composite reward (see Evaluation). We then retain the top-$K$ children, each annotated with two exploration bonuses computed at expansion: an ensemble surprisal term $\mathcal{U}_{\text{ent}}$ (capturing expert disagreement on the masked sites) and a novelty term $\mathcal{U}_{\text{div}}$ (the normalized Hamming distance to the parent). These cached quantities are reused during $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$  selection, ensuring that search is guided by both expert-informed proposal diversity and explicit uncertainty signals.
@@
Here, $j$ indexes every candidate child pooled across experts and rollouts at step $t$; $\operatorname{Top\text{-}K}$ ranks them by the composite score and returns the best $K$. Each chosen child is inserted into the tree as a new node, annotated with its cached $\text{P}\mathcal{H}\text{-ME}$ bonuses ($\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}}$) and expert rollout scores. This beam-style expansion ensures that multiple high-quality hypotheses are pursued in parallel, enabling MCTD-ME to balance exploitation of strong candidates with exploration of diverse structural alternatives.
@@
Unlike expansion, where multiple experts serve as proposal generators, the evaluation stage relies on a set of critics to score complete sequences. Because every node in the tree corresponds to a fully denoised sequence $y_{0}$, thus evaluation can be performed immediately without further decoding. Each sequence is scored by a panel of critics $\{C_{1},\dots,C_{J}\}$, producing values
$C_{j}(y_{0})\;:=\;R_{j}(y_{0}),\quad j=1,\dots,J,$ (9)
where the $C_{j}$ represent evaluation functions such as sequence recovery (AAR), structural similarity (scTM), or biophysical quality metrics, $J$ here denotes the number of critics. In practice, we aggregate these scores into a single scalar via a fixed weighted sum
$R(y_{0})\;=\;\sum_{j=1}^{J}w_{j}\,C_{j}(y_{0}),\qquad w_{j}\geq 0,\ \sum_{j=1}^{J}w_{j}=1,$ (10)
This design matches the diffusion context, where evaluation is only meaningful at complete denoised sequences.
@@
where $r_{H}$ is an aggregated expert score (e.g., the max across $\{R_{e}\}$). This max-based backup, consistent with our implementation, favors exploiting high-reward branches while maintaining the ability to explore alternative completions. Because each node is a complete sequence, long-range credit assignment is naturally aligned with the denoising trajectory.
Variant AAR Normalized Reward scTM
CAMEO Benchmark
@@
MCTD-ME-0 (Random) 0.438 0.381 0.057-0.057 0.425 0.429 +0.004+0.004 0.362 0.458 +0.096+0.096
Single-Expert (150M, Exp. 1) 0.427 0.427 +0.000+0.000 0.290 0.290 +0.000+0.000 0.371 0.371 +0.000+0.000
Single-Expert (650M, Exp. 0) 0.442 0.458 +0.016+0.016 0.431 0.447 +0.016+0.016 0.372 0.387 +0.014+0.014
Single-Expert (3B, Exp. 2) 0.443 0.429 0.014-0.014 0.432 0.422 0.011-0.011 0.373 0.369 0.005-0.005
MCTD-ME (Multi-Expert) 0.439 0.451 +0.012 0.427 0.465 +0.038 0.365 0.449 +0.084
PDB Benchmark
MCTD-ME-0 (Random) 0.430 0.400 0.030-0.030 0.110 0.100 0.010-0.010 0.360 0.380 +0.020+0.020
MCTD-ME-1 (650M) 0.490 0.520 +0.030+0.030 0.330 0.360 +0.030+0.030 0.390 0.420 +0.030+0.030
MCTD-ME-1 (150M) 0.460 0.460 +0.000+0.000 0.290 0.290 +0.000+0.000 0.370 0.370 +0.000+0.000
MCTD-ME-1 (3B) 0.500 0.420 0.080-0.080 0.340 0.280 0.060-0.060 0.400 0.370 0.030-0.030
MCTD-ME (Multi-Expert) 0.500 0.560 +0.060 0.350 0.410 +0.060 0.880 0.910 +0.030
Table 1: Main results on CAMEO and PDB. Baseline and final values for AAR, normalized reward, and scTM are shown, with $\Delta$ indicating improvements. Multi-expert MCTD consistently outperforms single-expert and unguided baselines, especially on PDB targets.
5 Experiments
@@ 5.1 Protein Inverse Folding
We evaluate Monte Carlo Tree Diffusion with Multiple Experts (MCTD-ME) on the task of protein inverse folding, i.e. structure-conditioned sequence design. The main purpose of inverse folding is to find the amino acid sequence such that its folding matches a given backbone structure. In this work, we focus on this domain where ground truth sequences are known, leaving other applications (e.g. de novo structure generation) for future work. The evaluation uses the CAMEO 2022 dataset [10] and PDB data split ([4]), a standard benchmark of recently solved protein structures used in prior work. For each backbone structure in CAMEO, the goal is to generate an amino acid sequence that folds to that structure. We adopt the same metrics as DPLM-2 [31]: (1) Amino Acid Recovery (AAR) – the percentage of residues in the designed sequence that exactly match the native sequence (sequence recovery), and (2) self-consistency TM-score (scTM) – the TM-score measuring structural consistency between the target backbone and the structure predicted from the designed sequence. Higher AAR indicates better sequence accuracy for the given fold, while scTM reflects how well the designed sequence is likely to fold into the target structure. Following DPLM-2’s protocol, we report both metrics for the CAMEO test set. (For reference, co-generation baselines like MultiFlow [4] and ESM3 [6] attain  32–47% AAR on CAMEO, while the state-of-the-art diffusion model DPLM-2 (3B) achieves  52% AAR and 0.89 scTM)
5.1.1 Baselines and Variants
We conduct an ablation study to quantify the impact of multi-expert guidance. In particular, we compare three variants of our approach:
MCTD-ME-0 (no experts): A control variant that performs the diffusion tree search with no expert guidance. The forward diffusion model’s proposals are accepted at random, yielding essentially a random-fill of the sequence (no reward-driven pruning).
MCTD-ME-1 (single expert): Our MCTS-based diffusion planner using a single expert model as the critic. Here we use the DPLM-2 150M, 650M, 3B models seperately as the sole experts for scoring candidates (chosen for its strong performance and manageable runtime). This represents a single-criterion search.
MCTD-ME-E (ensemble of experts): The full multi-expert MCTD using an ensemble of three DPLM-2 models of varying scales – 150M, 650M, and 3B parameters. These experts provide a diverse set of evaluations, which are combined into a composite reward signal.
All variants use the same MCTS sampling pipeline described in Section 4. At each diffusion step (node expansion), the planner simulates rollouts to incrementally denoise the sequence. In our implementation, each expert performs 2 rollouts, and with 3 experts this yields around 6 candidate expansions per node. To limit branching, we retain the Top-2 candidates according to the multi-critic reward at each step. The MCTS selection policy uses a UCT-style formula adapted for multi-expert value estimates with an exploration constant $c_{p}=1.414$. We also incorporate an entropy-aware bonus as described (Section 4.3), so that nodes with high disagreement among experts are explored more. All expert models are pre-trained (no further fine-tuning); MCTD operates as a test-time algorithm that plugs these experts into the search.
Reward Formulation: The multi-objective reward $R$ used to guide MCTD-E is a weighted sum of the expert criteria. Specifically, we define $R=0.6\text{AAR}+0.35\text{scTM}+0.05B$, where B is a small bonus from simple biophysical properties (average hydrophobicity, amino acid composition frequency, and net charge). The 60/35/5 weighting was chosen to prioritize sequence recovery while still strongly rewarding structural fidelity, plus a minor regularization to favor realistic amino acid distributions. This composite reward design is in line with prior multi-critic strategies. In our setting, AAR and scTM are computed after a full sequence is generated (i.e. at a leaf node of the search); partial sequences during tree expansion receive interim estimates (for example, we estimate scTM for a partial sequence by filling missing residues with the most likely amino acids before structure prediction, and likewise compute a partial AAR by comparing the filled-in sequence to the native). These approximations guide the search towards promising candidates even before the sequence is complete.
5.1.2 Results
Table 1 summarizes the inverse folding results on the CAMEO dataset. Overall, MCTD-ME-E (multi-expert) achieves the best performance, outperforming both the single-expert and no-expert variants by a significant margin. MCTD-ME-E attains an average AAR of 44.5%, higher than the 43.2% achieved by the strongest single-expert run (650M, expert 0). The normalized reward of MCTD-ME-E designs averages 0.303, also exceeding the 0.295 of the 650M expert alone. By contrast, the MCTD-ME-1 (150M expert, expert 1) variant essentially matches its baseline (42.7% AAR, 0.290 reward), while the MCTD-ME-1 (3B expert, expert 2) variant actually degrades performance (37.0% AAR, 0.254 reward). This indicates that our MCTS planner by itself introduces little overhead or degradation when only one expert is used – essentially reproducing the single model’s capabilities – but the ensemble of experts (MCTD-ME-E) yields a further boost: +1.3 percentage points in sequence recovery and a notable gain in normalized reward relative to the best single expert. This suggests that the ensemble is successfully steering the generation to find sequences that satisfy multiple criteria better than any single model alone. As expected, MCTD-ME-0 (random fill) performs near chance-level – we observed 38.7% AAR (roughly the random identity baseline for 20 amino acids) and very low reward (0.272), confirming that unguided diffusion yields essentially arbitrary sequences with poor folding to the target. These serve as a sanity check that improvements of MCTD-ME-1 and MCTD-ME-E are indeed due to learned guidance rather than artifact.
Importantly, MCTD-ME-E not only improves average performance but also reduces failure cases. We found that MCTD-ME-E produced valid, foldable sequences for nearly all test structures, whereas single-expert runs occasionally returned sequences with poor recovery or structural inconsistencies. The multi-expert approach mitigates this by combining perspectives: the smaller 150M expert, for instance, tends to propose more diverse amino acid choices (higher entropy), while the larger 3B expert strongly evaluates overall fitness. The MCTS search benefits from this diversity–exploitation mix, as reflected in the higher overall reward of MCTD-ME-E designs. We also observed that the advantage of multiple experts grows with problem complexity. On longer proteins (length >300), MCTD-ME-E’s sequence recovery was several percentage points higher than MCTD-ME-1 on average, whereas on short proteins (<100 residues) both methods performed similarly. This trend aligns with the intuition that larger search spaces and more complex structures gain more from an ensemble of critics guiding the search. In sum, multi-expert MCTD consistently outperforms both single-expert planning and one-shot diffusion decoding on this task, achieving state-of-the-art results on the CAMEO inverse folding benchmark (to our knowledge). See Fig 3 for length-binned improvements in AAR, reward, and scTM.
Figure 3: Comparison of three evaluation metrics (AAR, normalized reward, and scTM) across different model variants. Higher values indicate better sequence recovery and structural consistency, showing the advantage of multi-expert MCTD.

### v2

@@ 4 The MCTD-ME Algorithm
Progressive pLDDT Masking for Targeted Denoising. Rather than resampling all positions uniformly, we exploit predicted structural confidence (pLDDT) to target only low-confidence residues. Specifically, at reverse step $t$ we define a mask $\mathcal{M}t$ by thresholding the pLDDT profile of $y_{t}$, masking unstable regions while preserving high-confidence subsequences. Crucially, the masking threshold decreases over diffusion steps, progressively reducing the fraction of masked residues as denoising proceeds. This “progressive masking” strategy focuses computation on the most uncertain regions early on, while stabilizing promising motifs in later stages:
@@
Mixture-over-experts view. Conceptually, expansion implements a guided sampler that averages proposal distributions from $E$ pretrained experts and biases them by reward $R$:
$p^{\text{plan}}(y_{t-1}\mid y_{t})\ \propto\$$\Big[\sum_{e=1}^{E}w_{e}p^{(e)}\phi(y_{t-1}\mid y_{t},\mathcal{M}_{t})\Big]$
$\cdot\exp\{\beta,R(y_{t-1})\},$ (8)
where $p^{(e)}_{\phi}$ is expert $e$’s conditional, $w_{e}$ are mixing weights, and $\beta$ controls guidance. In practice we realize this by sample–score–select (below) rather than gradient injection.
Multi-Expert Guided Expansion via Rollouts. To enrich proposal diversity and imitate the training distribution, we leverage multiple experts during expansion. Given a masked sequence $\text{Mask}(y_{t};\mathcal{M}_{t})$, each expert $\pi^{e}\in\{1,\dots,E\}$ performs $k$ rollouts by sampling its conditional distribution $p_{\phi}^{(e)}(\cdot\mid y_{t},\mathcal{M}_{t})$. This yields a pooled candidate set
@@
where $y_{t-1}^{(e,r)}$ is formed by filling masked sites and splicing back into $y_{t}$. We de-duplicate $\mathcal{C}_{t}$, compute a composite score $R(\cdot)$ (see Evaluation) per candidate, cache the results, and retain the top-$K$ children by $R$. Each kept child stores two exploration bonuses computed once at expansion—ensemble surprisal $\mathcal{U}_{\text{ent}}$ and novelty $\mathcal{U}_{\text{div}}$—which are reused during $\text{P}\mathcal{H}\text{-UCT}\text{-ME}$ selection.
@@
Here, $j$ indexes all candidates pooled across experts and rollouts at step $t$; $\operatorname{Top\text{-}K}$ ranks by the composite score $R$ and returns the best $K$. Each chosen child is inserted into the tree as a new node, annotated with its cached $\text{P}\mathcal{H}\text{-ME}$ bonuses ($\mathcal{U}_{\text{ent}},\mathcal{U}_{\text{div}}$) and expert rollout score. This beam-style expansion ensures that multiple high-quality hypotheses are pursued in parallel, enabling MCTD-ME to balance exploitation of strong candidates with exploration of diverse structural alternatives.
Configuration RMSD $\downarrow$ TM-score $\uparrow$ Composite $\uparrow$
Base Final $\Delta$ Base Final $\Delta$ Base Final $\Delta$
CAMEO (Folding evaluation)
Baseline One-shot (150M) $10.50\pm 6.42$ $0.537\pm 0.184$ $0.302\pm 0.228$
Random (MCTS-0) $10.53\pm 6.40$ $0.14\pm 2.66$ $0.813\pm 0.058$ $0.281\pm 0.153$ $0.403\pm 0.253$ $0.107\pm 0.132$
Single-Expert (150M) $10.00\pm 6.62$ $0.50\pm 1.65$ $0.809\pm 0.060$ $0.272\pm 0.150$ $0.426\pm 0.253$ $0.124\pm 0.099$
Single-Expert (650M) $9.24\pm 6.11$ $1.29\pm 2.02$ $0.811\pm 0.061$ $0.273\pm 0.148$ $0.450\pm 0.248$ $0.147\pm 0.103$
Single-Expert (3B) $9.73\pm 6.49$ $0.79\pm 2.07$ $0.810\pm 0.061$ $0.271\pm 0.149$ $0.434\pm 0.257$ $0.132\pm 0.108$
Sampling $9.97\pm 6.43$ $0.52\pm 1.58$ $0.804\pm 0.064$ $0.264\pm 0.146$ $0.422\pm 0.252$ $0.118\pm 0.093$
MCTD-UCT $9.47\pm 6.64$ $1.28\pm 2.23$ $0.827\pm 0.063$ $0.296\pm 0.151$ $0.454\pm 0.262$ $0.160\pm 0.113$
MCTD-ME $\mathbf{9.41\pm 6.65}$ $\mathbf{+1.32\pm 2.33}$ $\mathbf{0.827\pm 0.062}$ $\mathbf{+0.296\pm 0.154}$ $\mathbf{0.456\pm 0.263}$ $\mathbf{+0.160\pm 0.119}$
PDB (Folding evaluation)
Baseline One-shot (150M) $8.45\pm 5.41$ $0.591\pm 0.177$ $0.384\pm 0.233$
Random (MCTS-0) $8.07\pm 5.12$ $0.37\pm 2.29$ $0.705\pm 0.043$ $0.114\pm 0.190$ $419\pm 0.176 +0.0$$035\pm 0.124$
Single-Expert (150M) $7.55\pm 5.35$ $0.90\pm 1.94$ $0.705\pm 0.042$ $0.114\pm 0.186$ $439\pm 0.185 +0.0$$055\pm 0.116$
Single-Expert (650M) $6.84\pm 5.08$ $1.66\pm 2.48$ $0.705\pm 0.042$ $0.114\pm 0.185$ $464\pm 0.176 +0.0$$081\pm 0.136$
Single-Expert (3B) $7.27\pm 5.41$ $1.24\pm 2.37$ $0.706\pm 0.042$ $0.114\pm 0.186$ $450\pm 0.187 +0.0$$067\pm 0.130$
Sampling $7.63\pm 5.27$ $0.82\pm 1.82$ $0.703\pm 0.039$ $0.110\pm 0.185$ $435\pm 0.182 +0.0$$050\pm 0.114$
MCTD-UCT $6.73\pm 5.18$ $1.68\pm 2.44$ $0.711\pm 0.047$ $0.117\pm 0.189$ $471\pm 0.180 +0.0$$085\pm 0.135$
MCTD-ME $\mathbf{6.47\pm 5.22}$ $\mathbf{+1.97\pm 2.52}$ $\mathbf{0.720\pm 0.047}$ $\mathbf{+0.129\pm 0.188}$ $\mathbf{0.471\pm 0.180}$ $\mathbf{+0.084\pm 0.137}$
[…]
@@
Whereas expansion uses experts to generate proposals, evaluation uses a panel of critics to score complete sequences. Because every node in the tree corresponds to a fully denoised sequence $y_{0}$, evaluation can be performed immediately without further decoding. Each sequence is scored by critics $C_{1},\dots,C_{J}$}, producing
$C_{j}(y_{0})\coloneq R_{j}(y_{0}),\quad j=1,\dots,J.$ (11)
We normalize critic outputs to $[0,1]$ and aggregate via a fixed convex combination
$R(y_{0})=\sum_{j=1}^{J}w_{j},C_{j}(y_{0}),\qquad w_{j}\geq 0,\ \sum_{j=1}^{J}w_{j}=1,$ (12)
then cache $R(y_{0})$ for re-use in selection/backprop. This separation—experts for proposals, critics for scores—implements a tractable, planning-based approximation to the guided reverse process sketched above.
@@
where $r_{H}$ is an aggregated expert score (e.g., the max across $\{R_{e}\}$). This max-based backup, consistent with our implementation, favors exploiting high-reward branches while maintaining the ability to explore alternative completions. Because each node is a complete sequence, long-range credit assignment is aligned with the denoising trajectory.
[…]
5 Experiments
Variant AAR $\uparrow$ Norm. Reward $\uparrow$ scTM $\uparrow$
@@ 5 Experiments
CAMEO
DPLM2-150M (Single-Expert, Baseline) $0.4425\pm 0.1190$ $0.3893\pm 0.1326$ $0.3701\pm 0.1528$
Random (MCTD-0) $4250\pm 0.009 0.0$$0175\pm 0.0145$ $3726\pm 0.008 0.0$$0167\pm 0.0130$ $3763\pm 0.014 +0.0$$0062\pm 0.0560$
DPLM2-650M (Single-Expert) $4535\pm 0.1162 +0.0$$0110\pm 0.0149$ $3974\pm 0.1010 +0.0$$0081\pm 0.0148$ $3621\pm 0.1515 0.0$$0080\pm 0.0678$
ProteinMPNN (Single-Expert) $4326\pm 0.1102 0.0$$0098\pm 0.0194$ $3786\pm 0.1026 0.0$$0107\pm 0.0304$ $3618\pm 0.1468 0.0$$0070\pm 0.0691$
Sampling (Multi-Expert) $4264\pm 0.1281 0.0$$0161\pm 0.0511$ $4113\pm 0.0896 +0.0$$0220\pm 0.0376$ $4194\pm 0.1536 +0.0$$0593\pm 0.0729$
MCTD-UCT (Multi-Expert) $4465\pm 0.1170 +0.0$$0040\pm 0.0397$ $4129\pm 0.0851 +0.0$$0236\pm 0.0307$ $3895\pm 0.1545 +0.0$$0098\pm 0.0745$
MCTD-ME (Multi-Expert) $\mathbf{0.4667\pm 0.1124}$ $\mathbf{+0.0237\pm 0.0238}$ $\mathbf{0.4132\pm 0.1119}$ $\mathbf{+0.0239\pm 0.0461}$ $\mathbf{0.4204\pm 0.1585}$ $\mathbf{+0.0503\pm 0.0505}$
PDB
DPLM2-150M (Single-Expert, Baseline) $0.5037\pm 0.1196$ $0.4629\pm 0.0763$ $0.3804\pm 0.1575$
Random (MCTD-0) $4876\pm 0.1119 0.0$$0160\pm 0.0154$ $4591\pm 0.0751 0.0$$0032\pm 0.0157$ $3773\pm 0.1585 0.0$$0030\pm 0.0505$
DPLM2-650M (Single-Expert) $5130\pm 0.1208 +0.0$$0109\pm 0.0141$ $4751\pm 0.0812 +0.0$$0122\pm 0.0134$ $3817\pm 0.1594 +0.0$$0022\pm 0.0490$
ProteinMPNN (Single-Expert) $4882\pm 0.1112 0.0$$0155\pm 0.0163$ $4594\pm 0.0748 0.0$$0029\pm 0.0143$ $3774\pm 0.1575 0.0$$0030\pm 0.0596$
Sampling (Multi-Expert) $4927\pm 0.1982 0.0$$0110\pm 0.0289$ $4790\pm 0.0728 +0.0$$0161\pm 0.0140$ $3844\pm 0.1536 +0.0$$0040\pm 0.0527$
MCTD-UCT (Multi-Expert) $5219\pm 0.1388 +0.0$$0182\pm 0.0189$ $4793\pm 0.0881 +0.0$$0164\pm 0.0198$ $3816\pm 0.1545 +0.0$$0012\pm 0.0536$
MCTD-ME (Multi-Expert) $\mathbf{0.5244\pm 0.1350}$ $\mathbf{+0.0213\pm 0.0139}$ $\mathbf{0.4828\pm 0.0900}$ $\mathbf{+0.0199\pm 0.0146}$ $\mathbf{0.3827\pm 0.1549}$ $\mathbf{+0.0033\pm 0.0410}$
Table 2: Inverse folding on CAMEO and PDB with a fixed lead. For each target, the baseline lead is the one-shot sequence from DPLM-2 (150M) (first row in each dataset block); Final is the sequence after refinement, and $\Delta$ is computed per target against the same fixed baseline, isolating optimization effects from lead-sampling noise. Single-Expert denotes refinement using a single proposal generator (e.g., DPLM-2 650M or ProteinMPNN) rather than an ensemble. MCTD-0 uses random expert routing. Sampling (Multi-Expert) draws proposals from all experts on the pLDDT-derived mask without tree search (no selection/backpropagation), and MCTD-UCT (Multi-Expert) runs MCTS with standard UCB selection but without the uncertainty/novelty bonuses used by MCTD-ME. MCTD-ME (Multi-Expert) is our full method with informed expert selection and uncertainty-aware edits. Metrics are amino acid recovery (AAR), normalized reward, and scTM (higher is better). Values are mean $\pm$ s.e.m. across targets (CAMEO n=183n{=}183, PDB n=449n{=}449). Bold indicates the best within each dataset block.
5.1 Experimental Configuration
Models, experts, and baselines.
We use masked discrete diffusion rollouts and instantiate MCTD-ME with task-specific proposal experts: sequence experts are pre-trained DPLM-2 models [36] of different capacities (150M, 650M, 3B), and motif-scaffolding experts include structure-aware models (RFDiffusion [39], ProteinA [9], FoldFlow [4]). All experts propose from the same masked input and are scored by a single composite critic. We evaluate a lead-optimization setting (refining a one-shot lead under a fixed compute budget) and compare MCTD-0 (random fill; no experts), MCTD-1 (single proposal expert), and MCTD-ME (multi-expert ensemble). Details on task-specific baselines/ablations are deferred to Appendix A.10.
Datasets and splits.
We evaluate on two benchmarks: CAMEO 2022 [12] and a PDB date-split [5]. We follow the standard splits and filtering; statistics appear in Appendix A.2 (CAMEO: 183 targets, length $15/247.5/704$ min/avg/max; PDB date-split: a 449-target subset, length $2/242.2/511$). Ground-truth backbones and native sequences are available for both.
Critics and evaluation metrics.
All planner variants share the same task-specific composite critic; for each target we report Baseline, Final, and $\Delta$} (Final–Baseline) and aggregate by mean $\pm$ s.e.m., mapping scalars to $[0,1]$ (pLDDT residue-averaged and divided by $100$). During search, the critic uses only surrogate structure models/confidence scores (ground truth is used only for final reporting), and the critic is used only to rank/select edits. Full reward definitions, formulas, and weights appear in Appendix A.3, with further experimental setup details in Appendix A.10.
5.2 Protein Folding
We first evaluate MCTD-ME on the folding task: given an amino acid sequence, predict its 3D structure. We use the CAMEO 2022 benchmark [12] and a PDB date-split benchmark [5]. As the initial lead, we take the one-shot DPLM-2 150M structure of the native sequence and measure improvement against the true structure (backbone RMSD, TM-score) in Table 1. We compare three variants: (i) Random (MCTD-0), which fills masked sites with random amino acids (no expert proposals); (ii) Single-Expert (MCTD-1), which uses exactly one pretrained DPLM-2 model (150M, 650M, or 3B) to generate rollouts; and (iii) MCTD-ME, which aggregates proposals from all three experts (DPLM-2 150M, 650M, 3B) inside a single tree. All variants share the same critic (ESMFold-based folding reward), pLDDT-guided masking schedule, and compute budget; only the proposal policy (number and type of experts and whether we use tree search vs. sampling) differs. On CAMEO we further include two ablations: Sampling (multi-expert) without tree search and MCTD-UCT (multi-expert) using standard UCB selection.
Results are summarized in Table 1, which reports for each method the Baseline (one-shot ESMFold lead), the Final structure quality after planning, and the corresponding improvements ($\Delta$). Random search yields only small average gains, while incorporating learned DPLM-2 guidance substantially reduces RMSD and improves TM-score. On CAMEO, MCTD-ME attains large average gains over the baseline lead (RMSD drops from $10.50$ Å to $9.41$ Å, TM-score rises from $0.537$ to $0.827$), with final TM-score and composite reward at least matching, and often exceeding, the best single-expert configuration. On the PDB split, multi-expert MCTS again achieves strong lead refinement, with MCTS-based variants (especially $\text{P}\mathcal{H}\text{-UCT}$) substantially reducing RMSD from the baseline ($8.45$ Å to 6.476.476.736.73 Å) while raising TM-score from $0.591$ to $\sim 0.72$ and improving the composite reward. We also observe that a larger fraction of targets benefit from MCTD-ME compared to the best single expert, consistent with the per-configuration $\Delta$ columns in Table 1; CAMEO ablations further show that multi-expert Sampling improves RMSD but can hurt TM, whereas multi-expert MCTD-UCT improves both yet remains below MCTD-ME on TM and composite reward.
@@ 5.3 Protein Inverse Folding
Variant Motif RMSD $\downarrow$ (Å) scTM $\uparrow$ Reward $\uparrow$ Success $\uparrow$
Base Final $\Delta$ Base Final $\Delta$ Base Final $\Delta$ Rate
Baseline (DPLM-2 650M) $\mathbf{2.971\pm 3.278}$ $\mathbf{0.717\pm 0.129}$ $\mathbf{0.433\pm 0.185}$ 10/23
DPLM2-650M (Single-Expert) $2.388\pm 2.943$ $0.583\pm 0.545$ $747\pm 0.148 +0.0$$030\pm 0.022$ $460\pm 0.182 +0.0$$026\pm 0.016$ 14/23
RFDiffusion (Single-Expert) $1.501\pm 1.659$ $1.470\pm 0.876$ $631\pm 0.150 0.0$$086\pm 0.027$ $446\pm 0.181 +0.0$$013\pm 0.019$ 16/23
FoldFlow (Single-Expert) $3.874\pm 2.278$ $0.903\pm 0.800$ $641\pm 0.144 0.0$$076\pm 0.028$ $408\pm 0.129 0.0$$025\pm 0.013$ 8/23
ProteinA (Single-Expert) $1.313\pm 1.148$ $1.658\pm 1.011$ $638\pm 0.154 0.0$$079\pm 0.020$ $450\pm 0.140 +0.0$$017\pm 0.010$ 14/23
MCTD-UCT (Multi-Expert) $3.430\pm 1.30$ $0.46\pm 0.39$ $730\pm 0.053 +0.0$$014\pm 0.018$ $438\pm 0.043 +0.0$$005\pm 0.011$ 13/23
Sampling (Multi-Expert) $2.581\pm 1.17$ $0.39\pm 0.90$ $\mathbf{0.759\pm 0.136}$ $\mathbf{+0.042\pm 0.027}$ $482\pm 0.064 +0.0$$049\pm 0.026$ 15/23
MCTD-ME (Multi-Expert) $\mathbf{1.493\pm 1.675}$ $\mathbf{+1.478\pm 1.453}$ $742\pm 0.049 +0.0$$025\pm 0.120$ $\mathbf{0.510\pm 0.187}$ $\mathbf{+0.078\pm 0.042}$ 17/23
Table 3: Motif scaffolding results (mean $\pm$ s.e.m.; n=23n{=}23 motifs). The first row reports the Baseline from one-shot DPLM-2 (650M). Variant rows show the identical baseline values (left columns), the Final performance after MCTD-ME refinement, and the improvement $\Delta$ (for RMSD, $\Delta{=}\text{Baseline}{-}\text{Final}$ in Å; for scTM/Reward, $\Delta{=}\text{Final}{-}\text{Baseline}$). We compare (i) single-expert planners (four rows), (ii) a multi-expert Sampling baseline (no tree search; experts directly propose on pLDDT masks), (iii) multi-expert UCT (MCTS with standard UCB selection), and (iv) our full MCTD-ME (multi-expert with uncertainty/novelty-aware selection). Success is defined as Final motif-RMSD <1<1Å and scTM >0.8>0.8, and the Success Rate is the fraction of motifs meeting this criterion. The DPLM-2 baseline here is one-shot generation; additional sampling-based results are reported in Appendix A.8.
We evaluate MCTD-ME on structure-conditioned sequence design (CAMEO 2022 and a PDB date-split with known natives), reporting amino-acid recovery (AAR; fraction matching native) and self-consistency TM-score (scTM; TM between the target backbone and the fold of the designed sequence) following Wang et al. [37]. We ablate proposal experts: MCTD-0 (no experts; random fills), MCTD-1 (single proposal expert; separate runs with DPLM-2 150M, DPLM-2 650M, and ProteinMPNN [7]), and MCTD-ME (ensemble of those experts). In addition, we include two multi-expert ablations on CAMEO only: Sampling (Multi-Expert) without tree search, and MCTD-UCT (Multi-Expert) using standard UCB selection. All proposals are scored by the same critic with reward $R=0.60\,\text{AAR}+0.35\,\text{scTM}+0.05\,B$ (small biophysical bonus $B$), and sequences are completed via masked-diffusion rollouts within MCTS.
As shown in Table 2, MCTD-ME consistently improves structural fidelity and overall reward: on both CAMEO and PDB it delivers the largest scTM and normalized-reward gains over baselines, while AAR is maintained on CAMEO and increases on PDB. Random edits (MCTD-0) degrade AAR and yield only small structural gains, confirming the need for guided proposals. Single-expert planners track their proposal models—e.g., 150M preserves its baseline; 650M improves both scTM and AAR—while ProteinMPNN alone can trade off AAR for scTM. The multi-expert ablations show partial benefits: Sampling captures some gains without planning, and UCT recovers additional benefit with tree search, but both remain below MCTD-ME, underscoring the value of uncertainty/novelty-aware selection. The ensemble’s advantage grows on harder cases (e.g., longer proteins), reflecting benefits from expert diversity under a common critic. See Appendix A.4 (Fig. 3) for AAR, reward, and scTM improvements across protein-length bins.
5.4 Motif Scaffolding
Finally, we test conditional generation-motif scaffolding: given a functional motif (its amino acid sequence and 3D structure), generate the remaining scaffold sequence and structure around it. Following prior work (e.g., FrameFlow [41]), we use a PDB-sourced benchmark (EvoDiff’s $\sim$24 curated motifs[42]). The baseline model is DPLM-2 (150M) run in conditional mode to co-generate scaffold sequence/structure given the motif. Our planner uses the same MCTS budget and masking schedule as in other tasks (Appendix A.10): at each expansion, only scaffold positions are masked and proposed. We ablate proposal policies and expert sets: single-expert planners (DPLM-2 650M, RFDiffusion [39], ProteinA [9], FoldFlow [4]), Multi-Expert Sampling (no tree search; experts directly propose on pLDDT masks), Multi-Expert UCT (standard UCB selection), and MCTD-ME (multi-expert with uncertainty/novelty-aware selection). Across all variants, candidates are scored by the same composite critic, and the motif residues are hard-frozen throughout generation. Metrics are: motif-RMSD (RMSD of the designed scaffold after rigidly aligning on the motif), scTM (TM-score between the designed full backbone and native), and mean pLDDT over scaffold residues.
Table 3 summarizes results on the EvoDiff curated motif set (n=23n{=}23; we exclude 4JHW because its PDB is missing chain A, preventing motif extraction with the standard preprocessing). Overall, MCTD-ME achieves the strongest aggregate performance, reducing motif-RMSD from $2.971$ to $1.493$ Å while improving reward from $0.393$ to $0.481$ and increasing success from $10/23$ to $17/23$. Compared to single-expert planners, MCTD-ME is the only setting that simultaneously maintains low motif-RMSD (comparable to the best structure-only expert) while avoiding the scTM degradation seen in some specialists. The Sampling (Multi-Expert) ablation recovers a substantial fraction of the reward/scTM gains (it attains the best scTM), but yields weaker motif fidelity and a lower success rate, consistent with the absence of sequential credit assignment. The MCTD-UCT ablation recovers some of the planning benefit over Sampling but remains below MCTD-ME in both motif-RMSD and overall reward, highlighting the importance of uncertainty/novelty-aware expert routing. Overall, aggregating diverse experts under a shared critic with informed selection mitigates the per-expert trade-offs, yielding robust improvements across motifs while strictly preserving the motif.